40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 14 00:06:09 2020, gautam, Update, LSC, Locking with POX for CARM 
    Reply  Tue Jul 14 01:55:03 2020, Koji, Update, LSC, Locking with POX for CARM 
    Reply  Tue Jul 14 17:28:29 2020, gautam, Update, LSC, Locking with POX for CARM PRFPMI_2020712sensMat.pdfLSCerrSigs.pdf
Message ID: 15476     Entry time: Tue Jul 14 00:06:09 2020     Reply to this: 15477   15481
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking with POX for CARM 

I tried using the POX_I error signal for the DC CARM_B path today a couple of times. Got to a point where the AO path could be engaged and the arm powers stabilized somewhat, but I couldn't turn the CARM_A path off without blowing the lock. Now the IMC has entered a temperemental state, so I'm abandoning efforts for tonight, but things to try tomorrow are:

  1. Check that the demod phase is set correctly
  2. With the CARM_B path engaged, measure some CARM OLTFs. Tonight, I was a bit over-optimistic I think, by expecting the scripted transition to take me all the way, but I think I'll have to fiddle around with the gains a bit.
  3. Check for offsets. The AO path should be AC coupled, but maybe the POX signal has some offset?

I have some data from a couple of days ago when the PRFPMI was locked as usual (CARM_B on REFL for both DC and AO paths), and the sensing lines were on, so I can measure the relative strength of the sensing lines in POX/REFL and get an estimate of what the correct digital gain should be.

The motivation here is to see if the sensing matrix looks any different with a modified locking scheme.

ELOG V3.1.3-